Help please
Find two consecutive integers that have a product of 20

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

Try 1*2
2*3
3*4
etc

Answer by mananth(16946)

let one integer be x
the other will be x+1
..
x(x+1)=20
x^2+x=20
-20
x^2+x-20=0
x^2+5x-4x-20=0
x(x+5)-4(x+5)=0
(x+5)(x-4)=0
x= -5 OR 4
-5+1=-4
(-5,-4)
...
..
4+1=5
(4,5)